OpenStudy (anonymous):

What value of x makes the equation true? –(9 + (–17)) = –9 + x x =

OpenStudy (anonymous):

@tkhunny @Basher @Hotchellerae21

OpenStudy (tkhunny):

Have you considered simplifying the expressions and solving for the value? Use your best Order of Operations.

OpenStudy (anonymous):

i keep getting 8 and -8

OpenStudy (tkhunny):

Well, only one is correct. Are you more confident with one or the other? –(9 + (–17)) = –9 + x Parentheses –(9 - 17) = –9 + x –(-8) = –9 + x 8 = –9 + x Add 9 8+9 = 9–9 + x 17 = x Check it –(9 + (–17)) = -(9-17) = -(-8) = +8 –9 + x ==> -9+17 = +8 Done.

OpenStudy (anonymous):

it is 17 i took the assesment